# Fading into darkness: A weak mass ejection and low-efficiency fallback accompanying black hole formation in M31-2014-DS1

### v1

@@ Abstract
Stellar-mass black holes (BHs) can form from the near-complete collapse of massive stars, causing them to abruptly disappear. The star M31-2014-DS1 in the Andromeda galaxy was reported to exhibit such a disappearance between 2014 and 2022, with properties consistent with the failed explosion of a M⊙ yellow supergiant leading to the formation of a M⊙ BH. We present mid-infrared (MIR) observations of the remnant obtained with the James Webb Space Telescope (JWST) and X-ray observations from the Chandra X-ray Observatory in 2024. The JWST MIRI/NIRSpec data reveal an extremely red source, showing strong blueshifted absorption from molecular gas (CO, CO2, H2O, SO2) and deep silicate dust features. Modeling the dust continuum confirms continued bolometric fading of the central source to (–8% of the progenitor luminosity), surrounded by a dust shell spanning au. Modeling of the molecular gas indicates M⊙ of gas expanding at km s-1 near the inner edge of the dust shell. No X-ray source is detected down to a luminosity limit of erg s-1. We show that the panchromatic observations are explained by (i) a low-energy ( erg) ejection of the outer H-rich progenitor envelope and (ii) a fading central BH powered by inefficient (% in mass) accretion of loosely bound fallback material. The analysis robustly establishes the bolometric fading of M31-2014-DS1 and provides the first cohesive insights into BH formation via low-energy explosions and long-term fallback.
@@ I Introduction
Despite decades of theoretical work and widespread observational evidence, the formation channels of stellar-mass black holes (BHs) remain poorly constrained by direct observations. BHs are understood to represent the end states of massive stellar evolution, with theory predicting multiple pathways to their formation (Burrows et al., 2025; Janka, 2025). These include both successful terminal explosions – in which the progenitor star ejects most of its core and outer envelope in a core-collapse supernova (CCSN; Janka et al. 2007; Smartt 2009) – and failed explosions, where the star undergoes near-complete implosion with little accompanying mass ejection (Kochanek et al., 2008; Lovegrove and Woosley, 2013). While CCSNe are now routinely observed in wide-field transient surveys (e.g. Perley et al. 2020; Pessi et al. 2025), identifying failed SNe and their weak mass-loss signatures has proven far more challenging, leaving this channel largely unconstrained (Kochanek et al., 2008; Byrne and Fraser, 2022; Basinger et al., 2021). Due to their expected low explosion energies ( erg), the most dramatic observational signature of failed SNe has been through the identification of luminous supergiants in nearby galaxies that disappear without an observed CCSN (Adams et al., 2017; De et al., 2026).
@@
While the conclusions regarding NGC 6946-BH1 were limited by the sparse temporal coverage of this (relatively) distant star, De et al. (2026) presented a new example of this phenomenon in the fortuitously nearby Andromeda galaxy. As in the case of NGC 6946-BH1, the source M31-2014-DS1 was identified as a luminous M⊙ yellow supergiant (YSG) that disappeared in optical light between 2014 and 2018, with exquisite MIR coverage of the fading from the NEOWISE mission. By combining a panchromatic archival dataset, De et al. (2026) demonstrated that the fading amplitude and timescale was well explained by the hydrogen-depleted progenitor envelope – that was partially initially expelled and subsequently accreted during the collapse. They showed that this ejection and partial fallback model also explains the observed properties of NGC 6946-BH1 – with its optical outburst and fading timescale indicative of a M⊙ YSG progenitor star with a M⊙ H-rich envelope, as previously also suggested from archival photometry (Humphreys, 2019). No neutrinos have been detected from a targeted search towards M31-2014-DS1 (Nakanishi et al., 2025; Suwa et al., 2025).
@@ II.1 JWST Observations
We obtained observations of M31-2014-DS1 with JWST as part of our Director’s Discretionary Time Program (DD 6809; PI: De). Observations were carried out between UT 2024-11-28 03:13 and 2024-11-28 10:55 using the Mid-infrared Instrument (MIRI; Wright et al. 2023) and the Near-infrared Spectrograph (NIRSpec; Böker et al. 2023). The MIRI observations consisted of a spectrum covering m obtained with the Low Resolution Spectrometer (LRS) and broadband imaging observations in the F1280W, F1500W, F1800W, F2100W and F2550W filters. The NIRSpec observations consisted of a medium resolution spectrum obtained in the G140H, G235H and G395H gratings. Both the MIRI LRS and NIRSpec observations included an offset target acquisition due to the expected faintness of the source at m and the very crowded field in the near-infrared bands (De et al. 2026; their Figure 1).
@@ III.1 Comparison to remnants of dusty transients
The MIR SED of the M31-2014-DS1 exhibits some distinctive features – i) an extremely red slope, with the spectrum rising by in luminosity from m, ii) a forest of strong, narrow absorption lines between m in the NIRSpec data, iii) a broad, deep absorption trough between m in the MIRI LRS data, which we identify with silicate absorption and iv) a fading SED at longer wavelengths inferred from the MIRI broadband imaging observations. These observations confirm the assumption of De et al. (2026) of O-rich circumstellar chemistry surrounding the star (which was previously not directly constrained). These characteristics are similar in many ways to those of dusty remnants of other types of low-energy transients and generally to dusty gas in O-rich circumstellar environments. Figure 2 compares the observed SED of M31-2014-DS1 to previously published observations of the remnant of a stellar merger in M31 (M31-LRN-2015 Karambelkar et al. 2025), the dustiest known Galactic stellar merger (OGLE-2002-BLG-360; Steinmetz et al. 2025), and the remnant of the failed SN candidate NGC 6946-BH1 (Beasor et al. 2024; Kochanek et al. 2024).
@@
The MIR spectrum of M31-2014-DS1 is remarkably similar to these events in their general red color as well deep silicate absorption near m. The similarity is particularly relevant for the case of NGC 6946-BH1, which was suggested by De et al. (2026) to be a failed SN from a marginally more massive progenitor with a larger H-envelope. Since MIR spectroscopy of its remnant are not available, Beasor et al. (2024) suggested that the photometric excess at m in NGC 6946-BH1 could be explained by polycyclic aromatic hydrocarbon (PAH) features, while Kochanek et al. (2024) argued that it was associated with silicate absorption from dust in O-rich gas. A comparison of M31-2014-DS1 and NGC 6946-BH1 confirms that the MIR SED of NGC 6946-BH1 is consistent with deep silicate absorption expected from an O-rich dust obscured source and generally the circumstellar environments of massive stars (Verhoelst et al., 2009). We attribute the similarity of the remnants of both M31-2014-DS1 and NGC 6946-BH1 to those of the stellar mergers as the characteristic appearance of dust obscured stars in O-rich environments.
@@
Figure 2 also compares their bolometric luminosity evolution relative to progenitor values. For M31-2014-DS1, we adopt the published bolometric light curve through 2022 from De et al. (2026) and extend it with the JWST epoch by performing a trapezoidal integration of the observed spectrum. In contrast to the failed SN candidates, the stellar mergers unambiguously leave remnants that remain over-luminous with respect to their progenitors for decades after eruption. The failed SN candidates instead exhibit substantially lower relative luminosities, supporting their association with terminal disappearance. Although the MIR coverage of NGC,6946-BH1 is too sparse to construct a well-sampled bolometric light curve, the densely sampled NEOWISE data for M31-2014-DS1, combined with the new JWST observations, demonstrate continued fading to % of the progenitor luminosity.
@@
Beasor et al. (2024) note similarities between the IR-luminous remnant of NGC 6946-BH1 and that of the post-eruption fading from the Great Eruption of Eta Carinae (Smith et al., 2011) – suggesting a possible analog if the progenitor of NGC 6946-BH1 was in an elevated luminosity state prior to outburst. However, the exquisite archival data on M31-2014-DS1 allows us to confirm that i) the progenitor of M31-2014-DS1 has remained at nearly constant optical brightness since its earliest published reports from observations in 1980 - 1990 ( yrs before the disappearance; Monet et al. 2003; Magnier et al. 1992), and ii) unlike the Eta Carinae remnant that has maintained a near-constant bolometric luminosity at the progenitor level for decades post-eruption (Mehner et al., 2019), M31-2014-DS1 continues to fade in luminosity (Figure 2).
@@ III.2 Dust continuum modeling
M31-2014-DS1 exhibits a clear peak of emission between m as expected from warm dust surrounding the remnant. To quantify the properties of the associated dust, we model the observed SED using the radiative transfer code DUSTY (Ivezic and Elitzur, 1997). We assume a spherically symmetric dust shell surrounding a central source described by a blackbody spectrum, together with astronomical warm silicate dust grains (given the observed deep absorption at m) with a single characteristic grain size and a density profile. The resulting free parameters of the dust model include the bolometric flux (; or equivalently luminosity ) and effective temperature () of the central source, the optical depth of the dust shell at 0.55 m (), the dust temperature at the inner radius of the shell (), the thickness of the shell where is the inner radius and is the outer radius of the dust shell, and the grain size ().
@@
For a given set of parameters, DUSTY computes the emergent spectral energy distribution, which we fit to the observed JWST NIRSpec, MIRI LRS and MIRI imaging observations by minimization using the affine-invariant Markov Chain Monte Carlo (MCMC) sampler emcee (Foreman-Mackey et al., 2013). Uniform priors are adopted for all parameters. Convergence of the MCMC chains is verified by visual inspection and by requiring the integrated autocorrelation times to be much shorter than the chain lengths. We adopt the median of the parameter posterior distributions as the best-fit value and its 68% confidence interval as the uncertainty. The resulting best-fit parameters and their uncertainties are shown in Table 1 and the best-fit model is overlaid with the data in Figure 1. The inner and outer radius of the shell are derived from the DUSTY output parameters of the best-fit model (Ivezic et al., 1999) and shown in Figure 1.
@@
The very high optical depth of the dust shell () implies substantial obscuration of the central source at optical and near-infrared wavelengths, naturally explaining the non-detection of the source in the G140H band. While the inferred optical depth is higher than estimated from the 2022 photometry (; De et al. 2026), the corresponding bolometric luminosity is lower – indicating continued fading of the source. We follow the relations in De et al. (2026) to estimate the total dust mass
where cm2 g-1 is the average visual opacity per unit dust mass.
@@ III.3 Molecular gas modeling
Figure 3 shows the NIRSpec and MIRI LRS molecular absorption between m after dividing the data by the resulting continuum model. We assume a physical model of a gas slab consisting of CO (both isotopologues of 12CO and 13CO), CO2, H2O and SO2. We use the slabspec module of the spectools-ir package (Salyk, 2022) to compute the wavelength-dependent optical depth – sampled on the same wavelength grid and spectral resolution as the data – of a given molecule given its column density (), temperature () and turbulent velocity (). We perform a minimization of the resulting transmission model spectrum simultaneously on the NIRSpec and MIRI/LRS data using the emcee code assuming uniform priors. We also include the common systemic velocity of the gas () as a free parameter (Table 1). We caution that the low resolution of the MIRI data severely limits estimation of a continuum model due the presence of narrow features; therefore, the MIRI data is not as well fit by the model. However, the dominant molecules in that region (H2O, SO2) do not significantly affect our interpretation.
@@
The high spectral resolution of the NIRSpec data shows that while the turbulent velocities for the individual molecules is similar to those in AGB star envelopes (Cami, 2002), the source exhibits a clear systemic blueshift motion of km s-1 relative to the barycentric frame. While this value is close to the known blueshift of M31 (Zhang et al., 2024), M31-2014-DS1 lies kpc away from the center of M31 and is subject to the galaxy rotation curve. To compare the observed velocity to the expected motion of the star, we follow the approach of Drout et al. (2009) and Massey et al. (2009) for red/yellow supergiants in M31 and use the seminal work on the rotation curve of M31 (Rubin and Ford, 1970). We find the expected systemic motion of the star to be km s-1 where we conservatively adopt the observed maximum scatter in the radial velocity motions of red supergiants in M31 (Massey et al., 2009) as the systematic uncertainty.
Comparison of the observed velocity to that expected for the source therefore indicates a net blueshift km s-1, indicative of an expanding layer of gas approaching the observer222We do not attempt to derive exact values for this velocity since the true bulk velocity is likely affected by optical depth effects within the expanding slab.. Evidence for expanding gas is also seen in the red edges of the absorption profiles, where weak emission features are detected in multiple lines (Figure 3), although a detailed modeling including optical depth effects in an expanding atmosphere is beyond the scope of this paper. While evidence for mass ejection and partial fallback presented in De et al. (2026) was only indirectly inferred from the long-term bolometric luminosity decay (given the lack of an optical outburst), these observations confirm the presence of ejected material. We discuss its implications for the nature of the source in Section IV.

### v2

@@ Abstract
Stellar-mass black holes (BHs) can form from the near-complete collapse of massive stars, causing them to abruptly disappear. The star M31-2014-DS1 in the Andromeda galaxy was reported to exhibit such a disappearance between 2014 and 2022, with properties consistent with the failed explosion of a M⊙ yellow supergiant leading to the formation of a M⊙ BH. We present mid-infrared (MIR) observations of the remnant obtained with the James Webb Space Telescope (JWST) and X-ray observations from the Chandra X-ray Observatory in 2024. The JWST MIRI/NIRSpec data reveal an extremely red source, showing strong blueshifted absorption from molecular gas (CO, CO2, H2O, SO2) and deep silicate dust features. Modeling the dust continuum confirms continued bolometric fading of the central source to (–8% of the progenitor luminosity), surrounded by a dust shell spanning au. Modeling of the molecular gas indicates M⊙ of gas expanding at km s-1 near the inner edge of the dust shell. No X-ray source is detected down to a luminosity limit of erg s-1. We show that the panchromatic observations are explained by (i) a low-energy ( erg) ejection of the outer H-rich progenitor envelope and (ii) a fading central BH powered by inefficient (% in mass) accretion of loosely bound fallback material. The analysis robustly establishes the bolometric fading of M31-2014-DS1 following its yr archival record of constant optical/infrared brightness, and provides the first cohesive insights into BH formation via low-energy explosions and long-term fallback.
@@ I Introduction
Despite decades of theoretical work and widespread observational evidence, the formation channels of stellar-mass black holes (BHs) remain poorly constrained by direct observations. BHs are understood to represent the end states of massive stellar evolution, with theory predicting multiple pathways to their formation (Burrows et al., 2025; Janka, 2025). These include both successful terminal explosions – in which the progenitor star ejects most of its core and outer envelope in a core-collapse supernova (CCSN; Janka et al. 2007; Smartt 2009) – and failed explosions, where the star undergoes near-complete implosion with little accompanying mass ejection (Kochanek et al., 2008; Lovegrove and Woosley, 2013). While CCSNe are now routinely observed in wide-field transient surveys (e.g. Perley et al. 2020; Pessi et al. 2025), identifying failed SNe and their weak mass-loss signatures has proven far more challenging, leaving this channel largely unconstrained (Kochanek et al., 2008; Byrne and Fraser, 2022; Basinger et al., 2021). Due to their expected low explosion energies ( erg), the most dramatic observational signature of failed SNe has been through the identification of luminous supergiants in nearby galaxies that disappear without an observed CCSN (Adams et al., 2017; De et al., 2024).
@@
While the conclusions regarding NGC 6946-BH1 were limited by the sparse temporal coverage of this (relatively) distant star, (De et al. 2024; hereafter D26) presented a new example of this phenomenon in the fortuitously nearby Andromeda galaxy. As in the case of NGC 6946-BH1, the source M31-2014-DS1 was identified as a luminous M⊙ yellow supergiant (YSG) that disappeared in optical light between 2014 and 2018, with exquisite MIR coverage of the fading from the NEOWISE mission. By combining a panchromatic archival dataset, D26 demonstrated that the fading amplitude and timescale was well explained by the hydrogen-depleted progenitor envelope – that was partially initially expelled and subsequently accreted during the collapse. They showed that this ejection and partial fallback model also explains the observed properties of NGC 6946-BH1 – with its optical outburst and fading timescale indicative of a M⊙ YSG progenitor star with a M⊙ H-rich envelope, as previously also suggested from archival photometry (Humphreys, 2019). No neutrinos have been detected from a targeted search towards M31-2014-DS1 (Nakanishi et al., 2025; Suwa et al., 2025).
@@ II.1 JWST Observations
We obtained observations of M31-2014-DS1 with JWST as part of our Director’s Discretionary Time Program (DD 6809; PI: De). Observations were carried out between UT 2024-11-28 03:13 and 2024-11-28 10:55 using the Mid-infrared Instrument (MIRI; Wright et al. 2023) and the Near-infrared Spectrograph (NIRSpec; Böker et al. 2023). The MIRI observations consisted of a spectrum covering m obtained with the Low Resolution Spectrometer (LRS) and broadband imaging observations in the F1280W, F1500W, F1800W, F2100W and F2550W filters. The NIRSpec observations consisted of a medium resolution spectrum obtained in the G140H, G235H and G395H gratings. Both the MIRI LRS and NIRSpec observations included an offset target acquisition due to the expected faintness of the source at m and the very crowded field in the near-infrared bands (D26; their Figure 1).
@@ III.1 Comparison to remnants of dusty transients
The MIR SED of the M31-2014-DS1 exhibits some distinctive features – i) an extremely red slope, with the spectrum rising by in luminosity from m, ii) a forest of strong, narrow absorption lines between m in the NIRSpec data, iii) a broad, deep absorption trough between m in the MIRI LRS data, which we identify with silicate absorption and iv) a fading SED at longer wavelengths inferred from the MIRI broadband imaging observations. The silicate absorption confirms the assumption of D26 of O-rich circumstellar chemistry surrounding the star (which was previously not directly constrained). These characteristics are similar in many ways to those of dusty remnants of other types of low-energy transients and generally to dusty gas in O-rich circumstellar environments. Figure 2 compares the observed SED of M31-2014-DS1 to previously published observations of the remnant of a stellar merger in M31 (M31-LRN-2015 Karambelkar et al. 2025), the dustiest known Galactic stellar merger (OGLE-2002-BLG-360; Steinmetz et al. 2025), and the remnant of the failed SN candidate NGC 6946-BH1 (Beasor et al. 2024; Kochanek et al. 2024).
@@
The MIR spectrum of M31-2014-DS1 is remarkably similar to these events in their general red color as well deep silicate absorption near m. The similarity is particularly relevant for the case of NGC 6946-BH1, which was suggested by D26 to be a failed SN from a marginally more massive progenitor with a larger H-envelope. Since MIR spectroscopy of its remnant are not available, Beasor et al. (2024) suggested that the photometric excess at m in NGC 6946-BH1 could be explained by polycyclic aromatic hydrocarbon (PAH) features, while Kochanek et al. (2024) argued that it was associated with silicate absorption from dust in O-rich gas. A comparison of M31-2014-DS1 and NGC 6946-BH1 suggests that the MIR SED of NGC 6946-BH1 is consistent with deep silicate absorption expected from an O-rich dust obscured source and generally the circumstellar environments of massive stars (Verhoelst et al., 2009). We attribute the similarity of the remnants of both M31-2014-DS1 and NGC 6946-BH1 to those of the stellar mergers as the characteristic appearance of dust obscured stars in O-rich environments.
@@
Figure 2 also compares their bolometric luminosity evolution relative to progenitor values. For M31-2014-DS1, we adopt the published bolometric light curve through 2022 from D26 and extend it with the JWST epoch by performing a trapezoidal integration of the observed spectrum. This pseudo-bolometric luminosity is computed by integrating the observed SED (Figure 1) between m, and performing a linear extrapolation to the edges of the wavelength range where the source falls below the detection threshold. In contrast to the failed SN candidates, the stellar mergers unambiguously leave remnants that remain over-luminous with respect to their progenitors for decades after eruption (Tylenda et al., 2013; Tylenda and Kamiński, 2016; Woodward et al., 2021). The failed SN candidates instead exhibit substantially lower relative luminosities, supporting their association with terminal disappearance. While Kashi and Soker (2017) have qualitatively suggested that the apparent remnant luminosity can be suppressed due to obscuration by a dusty disk, Kochanek (2024) have quantitatively shown that this suppression can be at most a factor of . Although the MIR coverage of NGC,6946-BH1 is sparse to construct a well-sampled bolometric light curve, the densely sampled NEOWISE data for M31-2014-DS1, combined with the new JWST observations, demonstrate continued fading to % of the progenitor luminosity.
There have been low-mass stellar mergers reported from optical searches in M31 in the last 40 years (Rich et al., 1989; Lipunov et al., 2017; Pastorello et al., 2021). Karambelkar et al. (2023) estimate that higher mass mergers (M⊙ primary stars, relevant to the progenitor of M31-2014-DS1) are rarer than lower mass objects. Therefore, the likelihood of detecting a detecting a failed SN (estimated % by D26) is no less than that for a high mass merger is M31 over the NEOWISE time baseline (% over yrs). This can be understood by noting that the fraction of massive stars estimated to undergo mergers (%; Sana et al. 2012) is very similar to the estimated fraction that undergo a failed SN (Neustadt et al., 2021).
@@
Beasor et al. (2024) note similarities between the IR-luminous remnant of NGC 6946-BH1 and that of the post-eruption fading from the Great Eruption of Eta Carinae (Smith et al., 2011) – suggesting a possible analog if the progenitor of NGC 6946-BH1 was in an elevated luminosity state prior (e.g. due to ongoing binary interaction) to outburst. However, the exquisite archival data on M31-2014-DS1 allows us to confirm that i) the progenitor of M31-2014-DS1 has remained at nearly constant optical/infrared brightness since its earliest published reports from observations in 1980 - 1990 ( mag for yrs before the disappearance; Monet et al. 2003; Magnier et al. 1992; Skrutskie et al. 2006), and ii) unlike the Eta Carinae remnant that has maintained a near-constant bolometric luminosity at the progenitor level for decades post-eruption (Mehner et al., 2019), M31-2014-DS1 continues to fade in luminosity (Figure 2).
@@ III.2 Dust continuum modeling
M31-2014-DS1 exhibits a clear peak of emission between m as expected from warm dust surrounding the remnant. To quantify the properties of the associated dust, we model the observed SED using the radiative transfer code DUSTY (Ivezic and Elitzur, 1997). We assume a spherically symmetric dust shell surrounding a central source described by a blackbody spectrum, together with warm astronomical silicate dust grains (given the observed deep absorption at m) with a single characteristic grain size and a density profile. The resulting free parameters of the dust model include the bolometric flux (; or equivalently luminosity ) and effective temperature () of the central source, the optical depth of the dust shell at 0.55 m (), the dust temperature at the inner radius of the shell (), the thickness of the shell where is the inner radius and is the outer radius of the dust shell, and the grain size ().
@@
For a given set of parameters, DUSTY computes the emergent spectral energy distribution, which we fit to the observed JWST NIRSpec, MIRI LRS and MIRI imaging observations by minimization using the affine-invariant Markov Chain Monte Carlo (MCMC) sampler emcee (Foreman-Mackey et al., 2013). Uniform priors are adopted for all parameters. Convergence of the MCMC chains is verified by visual inspection and by requiring the integrated autocorrelation times to be much shorter than the chain lengths. We adopt the median of the parameter posterior distributions as the best-fit value and its 68% confidence interval as the uncertainty. The resulting best-fit parameters and their uncertainties are shown in Table 1 and the best-fit model is overlaid with the data in Figure 1. The corresponding posterior distributions are shown in Appendix A. The inner and outer radius of the shell are derived from the DUSTY output parameters of the best-fit model (Ivezic et al., 1999) and shown in Figure 1.
@@
The very high optical depth of the dust shell () implies substantial obscuration of the central source at optical and near-infrared wavelengths, naturally explaining the non-detection of the source in the G140H band. While the inferred optical depth is higher than estimated from the 2022 photometry (; D26), the corresponding bolometric luminosity is lower – indicating continued fading of the source. The increasing optical depth results in a larger fraction of the total luminosity to emerge at long wavelengths; by directly integrating the best-fit DUSTY models, we estimate the fraction of the source luminosity emerging at m increases from % for the progenitor star to % and % for the 2022-2023 and 2025 remnant (observed with JWST), respectively. This indicates that only a small fraction of the total remnant luminosity is emerging at wavelengths longer than the JWST/MIRI coverage.
While the observed SED does not show evidence of a long-wavelength rising component that would indicate a cooler dust shell missed in the MIR observations, we can empirically constrain the presence of such a shell using the observed data. The outer dust temperature in our best-fit model is K; therefore we only investigate possible cooler shells with K. As the MIR SED and the effects of dust self-absorption (i.e. the silicate feature) is well described by the warm dust modeled here, we can assume that any outer cooler shell will be optically thin to its own emission. In such a case, the outer cooler SED can be described by a modified black-body function (Fox et al., 2010, 2011; Myers et al., 2024) using optical constants for Silicate dust (Draine and Li, 2001). Since a cold dust component would only appear at the longest wavelengths, we estimate the maximum luminosity allowed by scaling the component luminosity to match the observed F2550W flux222This is a very conservative maximum luminosity estimate since the inner hot dust shell also contributes to the luminosity at this wavelength.. For a K component, we find the maximum allowed luminosity to be % of the total measured remnant luminosity at MIR wavelengths. While a K component could still hide % of the remnant luminosity as the MIR SED, its implied radius ( au) would be unphysically large (implying ejecta velocities km s-1). We therefore rule out a cold dust component that could outshine the observed remnant MIR luminosity.
We follow the relations in D26 to estimate the total dust mass
where cm2 g-1 is the average visual opacity per unit dust mass. The measured dust mass is at the lower end of the distribution observed for core-collapse SNe at d (Wesson et al., 2015; Bevan et al., 2019); however, the measurement is very similar to the measured dust masses in the low-energy mass ejections observed for stellar mergers at similar phases (Karambelkar et al., 2025). As the ejected material is expected to be roughly of solar composition, such condensation is theoretically expected in low-energy transients due to grain condensation from metals (Bermúdez-Bustamante et al., 2024; González-Bolívar et al., 2024).
@@ III.3 Molecular gas modeling
Figure 3 shows the NIRSpec and MIRI LRS molecular absorption between m after dividing the data by the resulting continuum model. We assume a physical model of a gas slab consisting of CO (both isotopologues of 12CO and 13CO), CO2, H2O and SO2. We use the slabspec module of the spectools-ir package (Salyk, 2022) and the HITRAN (Rothman et al., 2010) database333https://hitran.org/hitemp/ to compute the wavelength-dependent optical depth (assuming local thermodynamic equilibrium) – sampled on the same wavelength grid and spectral resolution as the data – of a given molecule given its column density (), temperature () and velocity broadening (). We perform a minimization of the resulting transmission model spectrum simultaneously on the NIRSpec and MIRI/LRS data using the emcee code assuming uniform priors. We also include the common systemic velocity of the gas () as a free parameter (Table 1). We caution that the low resolution of the MIRI data severely limits estimation of a continuum model due the presence of narrow features; therefore, the MIRI data is not as well fit by the model. However, the dominant molecules in that region (H2O, SO2) do not significantly affect our interpretation.
@@
The high spectral resolution of the NIRSpec data shows that while the broadening velocities for the individual molecules is similar to those in AGB star envelopes (Cami, 2002), the source exhibits a clear systemic blueshift motion of km s-1 relative to the barycentric frame. While this value is close to the known blueshift of M31 (Zhang et al., 2024), M31-2014-DS1 lies kpc away from the center of M31 and is subject to the galaxy rotation curve. To compare the observed velocity to the expected motion of the star, we follow the approach of Drout et al. (2009) and Massey et al. (2009) for supergiants in M31 and use the seminal work on the rotation curve of M31 (Rubin and Ford, 1970). We find the expected systemic motion of the star to be km s-1 where we conservatively adopt the observed maximum scatter in the radial velocity motions of red supergiants (RSGs) in M31 (Massey et al., 2009) as the systematic uncertainty.
Comparison of the observed velocity to that expected for the source therefore indicates a net blueshift km s-1, indicative of an expanding layer of gas approaching the observer. Taking the maximum radial extent of the gas layer (out to au) together with the elapsed time since the start of the disappearance ( d; D26) provides a similar estimate444We do not attempt to derive exact values for this velocity since the true bulk velocity is likely affected by optical depth effects within the expanding slab. of km s-1. Evidence for expanding gas is also seen in the red edges of the absorption profiles, where weak emission features are detected in multiple lines (Figure 3), although a detailed modeling including optical depth effects in an expanding atmosphere is beyond the scope of this paper. While evidence for mass ejection and partial fallback presented in D26 was only indirectly inferred from the long-term bolometric luminosity decay (given the lack of an optical outburst), these observations confirm the presence of ejected material. We discuss its implications for the nature of the source in Section IV.